# Resource Allocation with Multi-Team Collaboration Based on Hamilton’s Rule

### v1

@@ IV Simulations
Figure 6: Simulation of Algorithm 1 with $m=4$ and $N=16$ while having all teams have identical mission evaluation functions with the same density function $\phi(q)$ in the locational cost in Equation 16 for all teams, but different team weights. (a) shows the initial random allocation of robots to teams, (b) shows the final robot allocation, and (c) shows the value of $\mathcal{G}$ at each collaboration iteration until collaboration is no longer possible (hence, getting the allocation in (b)).
@@ V Conclusion
Optimizing collaboration within multi-robot systems is critical for enhancing operational efficiency, scalability, and robustness in complex environments. In this paper, we presented an ecology-inspired framework for facilitating collaboration in multi-team environments that achieves optimal performance for the entire system through altruistic behavior. However, many open directions remain for this work, including incorporating time-varying importance and mission evaluation functions, heterogeneous agents, and other mission applications such as search and rescue and wildfire management.

### v2

@@ IV Simulations
Figure 6: Simulation of Algorithm 1 with $m=4$ and $N=16$ while having all teams have identical mission evaluation functions with the same density function $\phi(q)$ in the locational cost in Equation 12 for all teams, but different team weights. (a) The initial random allocation of robots to teams. (b) The final robot allocation. (c) The value of $\mathcal{G}$ at each collaboration iteration until collaboration is no longer possible (hence, getting the allocation in (b)).
@@ V Conclusion
Optimizing collaboration within multi-robot systems is critical for enhancing operational efficiency, scalability, and robustness in complex environments. In this paper, we presented an ecology-inspired framework for facilitating collaboration in multi-team environments that achieves optimal performance for the entire system through altruistic behavior, where robots are considered as shareable system resources. However, many open directions remain for this work, including incorporating time-varying importance and mission evaluation functions, heterogeneous agents, and other mission applications such as search and rescue and wildfire management.